OpenStudy (anonymous):

I have no idea where to begin can someone please help me out.... At 40 degrees north latitude the the number of hours of sunlight on a given day can be found by D(t)=12-3cos[2pie(t+10)/365] where t is measuredin days and t=0 is January 1st. 1. Find D'(t) 2. Find and interpret D'(59) (note t=59 corresponds to March 1st)

OpenStudy (anonymous):

This is the equation in the question\[D(t)=12-3\cos[\frac{ 2\pi(t+10) }{ 365}]\]